# Synth: Boosting Visual-Language Models with Synthetic Captions and Image Embeddings

### v1

@@ Abstract
The creation of high-quality human-labeled image-caption datasets presents a significant bottleneck in the development of Visual-Language Models (VLMs). We propose a novel approach that leverages the strengths of Large Language Models (LLMs) and image generation models to create synthetic image-text pairs for efficient and effective VLM training. Our method employs pretraining a text-to-image model to synthesize image embeddings starting from captions generated by an LLM. These synthetic pairs are then used to train a VLM. Extensive experiments demonstrate that the VLM trained with synthetic data exhibits comparable performance on image captioning, while requiring a fraction of the data used by models trained solely on human-annotated data. In particular, we outperform the baseline by 17% through augmentation with a synthetic dataset. Furthermore, we show that synthesizing in the image embedding space is 25% faster than in the pixel space. This research introduces a promising technique for generating large-scale, customizable image datasets, leading to enhanced VLM performance and wider applicability across various domains, all with improved data efficiency and resource utilization.
1 Introduction
@@
Visual-language models (VLMs) are quickly emerging as powerful tools for understanding visual and textual information. Their ability to combine these two modalities holds immense promise for application ranging from image captioning to visual question answering. While VLMs hold significant potential, their performance is often constrained by limited data availability. Recent breakthroughs demonstrate that pre-training VLMs on larger image-text pair datasets leads to significant improvements in downstream tasks (Li et al., 2022b; Hu et al., 2022). However, creating such datasets poses several challenges such as scarcity of paired data, high curation costs, low diversity and high imbalance in semantics, and the potentially noisy nature of data sourced from the internet (e.g., LAION (Schuhmann et al., 2021)). These factors often lead to laborious filtering and extended training times due to low signal-to-noise ratio, thus increasing overall resource consumption.
@@
This work tackles these limitations by introducing a novel approach that leverages pre-trained generative text and image models to efficiently create synthetic paired data for VLMs. Our approach uniquely generates both text and images synthetically, overcoming the reliance on real-world data and addressing the challenges of scarcity, cost, and noise.
| Generator | Method | Generator Model | Generator Range | Generated Pairs | Caption Class | Caption Type | Evaluation Setting |
| Mishra et al. (2022), | Mixed pairs such as | ||||||
| Greff et al. (2022), | (Segmentation, Images), | ||||||
| Simulation/Rendering Engines | Zheng et al. (2020), | Mix | Mix | (Optimal Flow, Videos), | N/A | N/A | Mix |
| de Melo et al. (2022) | (Depth Maps, Images) | ||||||
| Cascante-Bonilla et al. (2023) | Mix | Pixel | (Captions, Images) | Complex | Rule-based | Vision Encoder | |
| Azizi et al. (2023), | SD, | ||||||
| Off-the-shelf Image Generator | Fan et al. (2023) | Imagen, | Pixel | (Class, Images) | Simple | ImageNet Classes | Classifier |
| MUSE | |||||||
| Li et al. (2023c) | SD | Pixel | (Captions, Images) | Complex | Human Generated | VLM | |
| LLM Generated | |||||||
| Controlled Image Generator | Synth | MUSE | Embedding | (Captions, Images) | Complex | Human Generated | VLM |
@@
While synthetic data generation has been explored for various computer vision tasks such as image segmentation, optical flow prediction, or image classification (Mishra et al., 2022; Greff et al., 2022; Azizi et al., 2023; Fan et al., 2023; Li et al., 2023c), its application to both visual and textual modalities within VLM training is a significant advancement.
Moreover, instead of using off-the-shelf image generation models, we employ a controlled image generation approach. That is, we pre-trained our image generator on a dedicated dataset (also used for VLM training) to prevent unintended knowledge transfer from the image generator into the VLM. This strategy ensures a fair evaluation of the impact of synthetic data.
Finally, our framework can operate also at the embedding level; our text-to-image model generates image embeddings which feed directly into the VLM, bypassing pixel-space rendering. This paradigm shift dramatically reduces memory overhead and resource consumption while preserving the quality of synthetic training data.
@@
Demonstrated Effectiveness: Experiments will showcase significant performance gains in image captioning when using our synthetic data, assessed against real and synthetic-data baselines.
In conclusion, we offer insights into the future of VLM training, highlighting the potential for creating customized and potentially infinitely expandable datasets using text-driven approaches.
2 Related Works
VLMs
@@ 2 Related Works
Training From Synthetic Data
The generation of synthetic data for training machine learning models remains a highly active area of research. While numerous studies (Mishra et al., 2022; Cascante-Bonilla et al., 2023; Greff et al., 2022; Zheng et al., 2020; de Melo et al., 2022) explore the use of model-based rendering engines or simulators, the remarkable advancements in high-quality image generators have ignited a surge of interest in leveraging generative models for creating training data. This trend has significantly impacted a wide range of computer vision tasks, including semantic segmentation (Li et al., 2022a; Ros et al., 2016; Baranchuk et al., 2021; Tritrong et al., 2021; Li et al., 2021; Chen et al., 2019), human motion understanding (Varol et al., 2017; Ma et al., 2022; Guo et al., 2022), and more recently image classification (Azizi et al., 2023; Fan et al., 2023). Our work investigates the potential of data-driven generative models within the domain of visual-language models (VLMs). We focus on their application in downstream tasks such as image captioning, where the ability to process complex scene descriptions is important. From this aspect, our approach is closest to a concurrent work by Li et al. (2023c) where the goal is to replace faulty images in a captioning pipeline with their synthetic versions. However, our work distinguishes itself from Li et al. (2023c) by achieving strong performance while utilizing 40 times less paired data and only a quarter of the parameters. This demonstrates the potential for achieving both performance and efficiency through our proposed approach. Furthermore, unlike the mentioned works (Azizi et al. (2023); Fan et al. (2023); Li et al. (2023c), we avoid using an off-the-self image generator trained on larger datasets. Our work can also be likened to the concept of cycle consistency (Zhu et al., 2017) in visual-language models. This principle, where image-to-text and text-to-image conversion can be employed as a form of synthetic data training, albeit with extra gradients throughout the network, exploring image-text cycle consistency during sampling Li et al. (2023a) or training Li et al. (2023b) has been explored in recent works, demonstrating promising results.
Finally, we emphasize the efficiency of pipelines that seamlessly connect data generation and model training. While most prior work has concentrated on generating images in pixel space, we investigate the generation of image embeddings that can be directly integrated into the VLM. Our approach aligns with recent work in the scene graph classification domain (Sharifzadeh et al., 2021, 2022), which has explored synthetic scene graph embedding generation using pretrained object representations.
@@
3 Synth
Given the synthetic generation of text and then images, we refer to our method as Synth. Synth is a pipeline for training VLMs using generative models in addition to collected human data. In this section we introduce different components of this pipeline, namely Caption Generation, Image Generation and the full Synth model.
3.1 Caption Generation
@@
Make up a human-annotated description of an image that contains the following object: . The caption should be around 30-40 words long. Describing the different components of the scene in an objective and unbiased way. Do not add subjective judgments about the image, it should be as factual as possible. Do not use fluffy, poetic language. Respond only with the caption itself, beginning with “This is an image of”.
@@ 3.2 Image Generation
Off-the-shelf image generator models, trained on vast image-text datasets, possess an unfair advantage for studying the effect of synthetic images in VLM training. Their exposure to extensive human annotated data might also enable them to produce images highly reminiscent of the original training set. To ensure fair evaluation, we propose pre-training a text-to-image generator in a controlled environment. This generator is trained on a human annotated image-caption dataset that will also be utilized for VLM training, alongside the synthetic data. By doing so, we effectively eliminate the effects of external data on the VLM training.
Our architecture is similar to Chang et al. (2023) and was chosen for its superior inference efficiency due to its use of discrete image tokens and its parallel decoding capability, which dramatically reduces the number of sampling iterations compared to auto-regressive models. Additionally, MUSE leverages a transformer architecture, further enhancing efficiency (Chang et al., 2023). Figure 3 shows some examples of generated images compared with original ones.
@@ 3.2.1 Training
To pretrain our image generator given paired image and texts, we embed the texts with a pre-trained language model and the images with a pre-trained VQ-GAN (Esser et al., 2021) (refer to A.1.1 in the appendix for the details). The pre-trained language model is the same used in our VLM and it is a reduced version of Chinchilla (Hoffmann et al., 2022) that is using 400 million parameters. Following the approach in Chang et al. (2023), we apply a masking procedure with a cosine scheduling function. This function gradually increases the masking ratio during training, biasing the model towards learning from images with larger masked regions. On average, around 64% of the VQ tokens are replaced with a special “dropped token”. The noisy VQ tokens and the embedded text are then fed into a transformer model. The model’s objective is to predict the original, unmasked VQ tokens. The loss function is the cross-entropy between the predicted VQ tokens and the masked ones:
where denotes the VQ tokens computed from the ground-truth image , denotes the ground-truth caption, and denotes the indices of the masked tokens.
@@ 3.2.2 Inference
During inference, we use an iterative procedure to generate images. At first iteration, the predicted VQ tokens are all initialized with the “dropped token” representing the masked entries. At each upcoming iteration, the text embedding and the predicted VQ tokens until that stage are fed into the transformer and the model predicts the masked tokens. A cosine schedule determines which masked tokens have the highest prediction confidence during decoding. These tokens are unmasked, progressively reducing the masked token set. This process repeats for 24 iterations, gradually refining the generated VQ-tokens.
Additional details for our text-to-image generator model are reported in A.1.3 in the appendix.
3.3 Synth VLM architecture
As described in Section 3.2, our text-to-image model generates image embeddings (VQ tokens) through an iterative caption-conditioned denoising process. A VQ-GAN decoder could be used to convert these tokens into human-viewable images. However, we emphasize that the image embeddings themselves provide a compressed representation of an image, making the decoder step unnecessary for certain applications.
Therefore, to enhance efficiency for training the VLM from synthetic data, we design the VLM to enable the bypass of pixel-space processing. This is achieved by setting its vision encoder to be identical to the VQ-GAN backbone used for our image generator. This enables seamless interaction between the VLM and synthetically generated image embeddings and eliminates the computationally expensive decoding stage of VQ-GAN for the image generator and the encoding stage for the VLM when training on synthetic data. At the same time, it allows us to train from images in pixel space when using human annotated data. Overall, our VQ-based design removes the need for costly conversions to and from pixel space, resulting in significant time and disk space savings as will be shown in Section 4. To harness the efficiency of discrete tokens during training and sampling, while maintaining the rich image understanding provided by soft embeddings, we convert each discrete token to its soft embedding using the codebook before feeding them to the VLM.
On top of the VQ backbone in our VLM, we incorporate a Perceiver Resampler component (Jaegle et al., 2021) to cross-attend to the VQ tokens extracted from our backbone. Similar to Flamingo (Alayrac et al., 2022), we employ a frozen language model to cross-attend to the image representation and autoregressively generate text. While the self-attention layers in the language model remain frozen, the cross-attention layers between the language and vision components are trainable, allowing the model to learn effective multimodal representations.
@@ 3.3 Synth VLM architecture
where denotes the image, represents the l-th language token, and represents all tokens preceding .
As shown in Figure 4 by combining the components introduced in the previous parts, we can train VLMs either from human annotated data or synthetic image-captions pairs such that the VLM is conditioned on the synthetic image embedding with the cross-entropy loss being between the synthetic captions and the VLM’s predictions.
@@ 4.1.1 Datasets
For training we used the following datasets in our experiments:
| Method | Real | Synth | #Real Data | #Synth Data | CIDEr-COCO () | CIDEr-Flickr-30 () |
| Baseline | CCv2 | - | 10.1M | - | 22.1 | 11.3 |
| Synth | CCv2 | LTIP | 10.1M | 330M | 28.7 | 12.9 |
| CCv2 | WebLI | 10.1M | 350M | 27.6 | 12.3 | |
| Gold | CCv2+LTIP | - | 340.1M | - | 27.6 | 13.2 |
| CCv2+WebLI | - | 360.1M | - | 30.7 | 15.1 |
Conceptual Captions v2 (Changpinyo et al., 2021) is a dataset with 12 million image-text pairs that are collected by automatically extracting images and their corresponding “alt-text” descriptions from various websites. In our experiments we use Conceptual Captions v2 as the main source of human-annotated data whereas for the other datasets we might use synthetically generated images/texts depending on the experiment.
WebLI (Chen et al., 2022) has around 350 million image and alt text pairs with high text quality in English.
GenPair refers to 1 million fully synthetic captions that are created as described in Section 3.1, paired with synthetic images generated on-the-fly during the VLM training using our text-to-image generator.
For evaluation we used:
MS-COCO (Chen et al., 2015): we evaluate the performance of our models on the commonly used test set of MS-COCO under zero-shot and fine-tuning settings. For the finetuning settings we use the training set MS-COCO. We use the Karpathy split (Karpathy & Fei-Fei, 2015) for evaluation. We use CIDEr score (Vedantam et al., 2015) as our metric.
Flickr-30k (Plummer et al., 2015): again, we evaluate performances on the Karpathy split.

### v2

# Synth2: Boosting Visual-Language Models with Synthetic Captions and Image Embeddings

@@ Abstract
The creation of high-quality human-labeled image-caption datasets presents a significant bottleneck in the development of Visual-Language Models (VLMs). In this work, we investigate an approach that leverages the strengths of Large Language Models (LLMs) and image generation models to create synthetic image-text pairs for efficient and effective VLM training. Our method employs a pretrained text-to-image model to synthesize image embeddings from captions generated by an LLM. Despite the text-to-image model and VLM initially being trained on the same data, our approach leverages the image generator’s ability to create novel compositions, resulting in synthetic image embeddings that expand beyond the limitations of the original dataset. Extensive experiments demonstrate that our VLM, finetuned on synthetic data achieves comparable performance to models trained solely on human-annotated data, while requiring significantly less data. Furthermore, we perform a set of analyses on captions which reveals that semantic diversity and balance are key aspects for better downstream performance. Finally, we show that synthesizing images in the image embedding space is 25% faster than in the pixel space. We believe our work not only addresses a significant challenge in VLM training but also opens up promising avenues for the development of self-improving multi-modal models.
keywords:
synthetic data, visual language models, multimodal self-improvement, open-ended learning1 Introduction
@@
Visual-language models (VLMs) are quickly emerging as powerful tools for understanding visual and textual information. Their ability to combine these two modalities holds immense promise for applications ranging from image captioning to visual question answering. While VLMs hold significant potential, their performance is often constrained by limited data availability. Recent breakthroughs demonstrate that pre-training VLMs on larger image-text pair datasets leads to significant improvements in downstream tasks (Li et al., 2022b; Hu et al., 2022). However, creating such datasets poses several challenges such as scarcity of paired data, potentially noisy nature of data sourced from the internet (e.g., LAION (Schuhmann et al., 2021)), high curation costs, low diversity and high imbalance in semantics. These factors often lead to laborious filtering and extended training times due to low signal-to-noise ratio, thus increasing overall resource consumption (Figure 1A).
@@
This work tackles these limitations by introducing an efficient approach that leverages pre-trained generative text and image models to create synthetic paired data for VLMs (see Figure 1B). Our approach uniquely synthesizes both text and images, overcoming the reliance on real-world data and addressing the challenges of scarcity, cost, and noise.
Additionally, the proposed framework operates at both the pixel and embedding levels (see Figure 1C), enabling us to train the VLM from either real images or synthetic image embeddings. This bypasses the need for pixel-space rendering for both the image generator and the VLM encoder. This paradigm shift significantly reduces memory overhead and resource consumption while maintaining the quality of synthetic training data.
@@
While synthetic data generation has been explored for various computer vision tasks such as image segmentation, optical flow prediction, or image classification (Mishra et al., 2022; Greff et al., 2022; Azizi et al., 2023; Fan et al., 2023; Li et al., 2023c), its application to both visual and textual modalities within VLM training is a significant advancement. Furthermore, a potential pitfall in studying the impact of synthetic data in training new models lies in the inherent advantage of off-the-shelf image generators. These models, trained on massive image-text datasets, may inadvertently equip VLMs with knowledge already embedded in the image generators’ own pre-training data. This overlap could mask the true effectiveness of synthetic images, leading to improvements that might be equally achievable by simply training VLMs on the original image generator dataset.
To isolate the unique contribution of synthetic images, we propose a controlled study. Both the text-to-image generator and the VLM undergo pre-training on the exact same dataset. This alignment ensures that any subsequent improvements observed in the VLM after fine-tuning on synthetic data can be confidently attributed to the novel compositions and representations generated, rather than pre-existing knowledge inherited from the original dataset.
@@
Demonstrated Effectiveness: Experiments will showcase significant performance gains in 3 domains (scene description, scene understating QA and Extrernal Knowledge QA) when using our synthetic data, assessed against real and synthetic-data baselines.
In conclusion, we offer insights into the future of VLM training, highlighting the potential for creating customized datasets, open-ended training and self-improving multimodal models.
2 Related Work
Generator | Method | Generator Model | Generated Set | Caption Class | Caption Type | Evaluation Setting | |
Canonical Concept Mapping | Sharifzadeh et al. (2021) | Linear | Scene Graph Embedding | Scene Graphs | |||
Sharifzadeh et al. (2022) | Complex Text | Human Generated | SG Classification | ||||
Mishra et al. (2022) | Mixed pairs such as | ||||||
Greff et al. (2022) | (Segmentation, Images) | ||||||
Simulation/Rendering Engine | Zheng et al. (2020) | Mix | (Optimal Flow, Videos) | N/A | N/A | Mix | |
de Melo et al. (2022) | (Depth Maps, Images) | ||||||
Cascante-Bonilla et al. (2023) | Mix | (Captions, Images) | Complex Text | Rule-based | Vision Encoder | ||
Azizi et al. (2023) | SD | ||||||
Off-the-shelf Image Generator | Fan et al. (2023) | Imagen | Images | Single Word | ImageNet Classes | Classifier | |
MUSE | |||||||
Li et al. (2023c) | SD | Images | Complex Text | Human Generated | VLM | ||
Controlled Image Generator | Synth2 | MUSE | Text, Embeddings & Images | Complex Text | Human & LLM Generated | VLM |
@@ 2 Related Work
Generation of synthetic data for training machine learning models remains a highly active area of research. While numerous studies (Mishra et al., 2022; Cascante-Bonilla et al., 2023; Greff et al., 2022; Zheng et al., 2020; de Melo et al., 2022) explore the use of model-based rendering engines or simulators, the remarkable advancements in high-quality image generators have ignited a surge of interest in leveraging generative models for creating training data. This trend has significantly impacted a wide range of computer vision tasks, including semantic segmentation (Li et al., 2022a; Ros et al., 2016; Baranchuk et al., 2021; Tritrong et al., 2021; Li et al., 2021; Chen et al., 2019), human motion understanding (Varol et al., 2017; Ma et al., 2022; Guo et al., 2022), and more recently image classification (Azizi et al., 2023; Fan et al., 2023). Our work investigates the potential of data-driven generative models within the domain of visual-language models (VLMs). We focus on their application in downstream tasks, where the ability to process complex scene is important. From this aspect, our approach is closest to a concurrent work by Li et al. (2023c) where the goal is to replace faulty images in a captioning pipeline with their synthetic versions. However, our work distinguishes itself from Li et al. (2023c) by achieving strong performance while utilizing 40 times less paired data and only a quarter of the parameters. This demonstrates the potential for achieving both performance and efficiency through our proposed approach. Furthermore, unlike the mentioned works (Azizi et al., 2023; Fan et al., 2023; Li et al., 2023c), we avoid using an off-the-self image generator trained on larger datasets. This approach prevented knowledge transfer into the VLM from a model trained on large amounts of paired data, which is critical as such knowledge transfer would have obscured our ability to scientifically assess the unique contribution of synthetic data.
Our work can also be likened to the concept of cycle consistency (Zhu et al., 2017) in visual-language models. This principle, where image-to-text and text-to-image conversion can be employed as a form of synthetic data training, albeit with extra gradients throughout the network, exploring image-text cycle consistency during sampling (Li et al., 2023a) or training (Li et al., 2023b) has been explored in recent works, demonstrating promising results.
Finally, we emphasize the efficiency of pipelines that seamlessly connect data generation and model training. While most prior work has concentrated on generating images in pixel space, we investigate the generation of image embeddings that can be directly integrated into the VLM. Our approach aligns with recent work in the scene graph classification domain (Sharifzadeh et al., 2021, 2022), which has explored synthetic scene graph embedding generation using pretrained canonical class representations.
@@
3 Synth2
Given the synthetic generation of text and then images, we refer to our method as Synth2. Synth2 is a pipeline for training VLMs using generative models in addition to collected human data. In this section we introduce different components of this pipeline, namely Caption Generation, Image Generation and the full Synth2 model (see Figure 3 for details).
3.1 Synthetic Caption Generation
@@
Make up a human-annotated description of an image that contains the following object: . The caption should be around 30-40 words long. Describe the different components of the scene in an objective and unbiased way. Do not add subjective judgments about the image, it should be as factual as possible. Do not use fluffy, poetic language. Respond only with the caption itself, beginning with “This is an image of”.
@@ 3.2 Image Generation
Our image generator architecture is similar to Chang et al. (2023) and was chosen for its superior inference efficiency due to parallel decoding and its use of discrete image tokens that dramatically reduces the number of sampling iterations compared to auto-regressive models. Additionally, MUSE leverages a transformer architecture, further enhancing efficiency (Chang et al., 2023). As discussed in Section 1, to mitigate the potential for knowledge overlap between off-the-shelf image generators and VLMs, we deviate from standard practice. Instead of employing pre-trained image generators, we intentionally pre-train both our text-to-image generator and VLM on an identical dataset. This controlled setup ensures that any performance gains observed in the VLM after fine-tuning on the synthetic data are directly attributable to the unique contributions of the generated images, rather than inherited knowledge from the original dataset. Hence, we pretrain an image generator on a image-caption dataset that will also be utilized for VLM pretraining (i.e. Conceptual Captions v2 (Changpinyo et al., 2021) see Section 4).
@@ 3.2.1 Training
To pretrain our image generator given paired image and texts, we embed the texts with a pre-trained language model and the images with a pre-trained VQ-GAN (Esser et al., 2021) (refer to A.1.1 in the appendix for the details). The pre-trained language model is the same used in our VLM and it is a reduced version of Chinchilla (Hoffmann et al., 2022) using 400 million parameters. Following the approach in Chang et al. (2023), we apply a masking procedure with a cosine scheduling function. This function gradually increases the masking ratio during training, biasing the model towards learning from images with larger masked regions. On average, around 64% of the VQ tokens are replaced with a special “dropped token”. The noisy VQ tokens and the embedded text are then fed into a transformer model. The model’s objective is to predict the original, unmasked VQ tokens. The loss function is the cross-entropy between the predicted VQ tokens and the masked ones:
where denotes the VQ tokens computed from the ground-truth image , denotes the ground-truth caption, and denotes the indices of the masked tokens. Our text-to-image generator was trained only on 10.1 millions text-image pairs from Conceptual Captions V2 (Changpinyo et al., 2021) (see Section 4 for details).
@@ 3.2.2 Inference
During inference, we use an iterative procedure to generate images. At first iteration, the predicted VQ tokens are all initialized with the “dropped token” representing the masked entries. At each upcoming iteration, the text embedding and the predicted VQ tokens until that stage are fed into the transformer and the model predicts the masked tokens. A cosine schedule determines which masked tokens have the highest prediction confidence during decoding. These tokens are unmasked, progressively reducing the masked token set. This process repeats for 24 iterations, gradually refining the generated VQ-tokens. Additional details for our text-to-image generator model and examples of generated images, given ground truth annotations compared to the original ones, are reported in A.1.3 in the appendix.
3.3 Synth2 VLM architecture
As described in Section 3.2, our text-to-image model generates image embeddings through an iterative caption-conditioned denoising process. A VQ-GAN decoder could be used to convert these tokens into human-viewable images. However, we emphasize that the image embeddings themselves provide a compressed representation of an image, making the decoder step unnecessary for certain applications.
Therefore, to enhance efficiency for training the VLM from synthetic data, we design the VLM to enable the bypass of pixel-space processing. This is achieved by setting its vision encoder to be identical to the VQ-GAN backbone used for our image generator. This enables seamless interaction between the VLM and synthetically generated image embeddings and eliminates the computationally expensive decoding stage of VQ-GAN for the image generator and the encoding stage for the VLM when training on synthetic data. At the same time, it still allows us to train from images in pixel space when using human annotated data. This design feature is particularly important since in order to curtail model collapse effects (Shumailov et al., 2023), we need to mix synthetic and real data during finetuning (Gerstgrasser et al., 2024).
Overall, our VQ-based design removes the need for costly conversions to and from pixel space, resulting in significant time and disk space savings as will be shown in Section 4. To harness the efficiency of discrete tokens during training and sampling, while maintaining the rich image understanding provided by soft embeddings, we convert each discrete token to its soft embedding using the codebook before feeding them to the VLM.
On top of the VQ backbone in our VLM, we incorporate a Perceiver Resampler component (Jaegle et al., 2021) to cross-attend to the VQ tokens extracted from our backbone. Similar to Flamingo Alayrac et al. (2022), we employ a frozen language model to cross-attend to the image representation and autoregressively generate text. While the self-attention layers in the language model remain frozen, the cross-attention layers between the language and vision components are trainable, allowing the model to learn effective multimodal representations.
@@ 3.3 Synth2 VLM architecture
where denotes the image, represents the l-th language token, and represents all tokens preceding . As shown in Figure 3 by combining the components introduced in the previous parts, we can train VLMs either from human annotated data or synthetic image-captions pairs such that the VLM is conditioned on the synthetic image embedding with the cross-entropy loss being between the synthetic captions and the VLM’s predictions.
@@ 4.1.1 Datasets
For training we used four datasets: (1) Conceptual Captions v2 (Changpinyo et al., 2021) (CCv2) is a dataset with 12 million image-text pairs that are collected by automatically extracting images and their corresponding “alt-tex” descriptions from various websites. In our experiments we use CCv2 as the main source of human-annotated data whereas for the other datasets we might use synthetically generated images/texts depending on the experiment. (2) WebLI (Chen et al., 2022) has around 350 million image and alt text pairs with high text quality in English. (3) LTIP (Alayrac et al., 2022) has 312 million web collected image and text pairs. (4) GenPair refers to 1 million fully synthetic captions paired with synthetic images generated that we create on-the-fly during the VLM training using our text-to-image generator. The details are described in Section 3.1.
We tested our model on 3 different domains:
Scene description: for this task we used MS-COCO (Chen et al., 2015) and we evaluate the performance of our models on the commonly used test set of MS-COCO under zero-shot and fine-tuning settings. For the finetuning settings we use the training set MS-COCO. We use the Karpathy split (Karpathy and Fei-Fei, 2015) for evaluation. Another dataset we used for this class is Flickr-30k (Plummer et al., 2015): again, we evaluate performances on the Karpathy split. We use CIDEr score (Vedantam et al., 2015) as our metric.